OpenStudy (hba):

A certain account should be opened with an investment for a positive integer number of years.no other diposit or withdrawals are permitted and the account earns 7% intrest that is compounded anually.if $400 were invested in the account for x years,what is the smallest possible value of x such that at the end of x years,the amount in the account will be atleast 3 times the initial investment?? (sat sub.test math lvl2)

OpenStudy (hba):

@Wired

OpenStudy (anonymous):

a little less complex please!

OpenStudy (anonymous):

Haven't done this in years, but I THINK this is how you do it: http://math2.org/math/general/interest.htm \[P = C (1 + \frac{r}{n})^{nt}\] where P = future value C = initial deposit r = interest rate (expressed as a fraction: eg. 0.06) n = # of times per year interest in compounded t = number of years invested \[\Large 3*400 \le 400 (1 + \frac{0.07}{1})^{1*x}\]